In a survey, 11 people were asked how much they spent on their child's last birthday gift. The results were roughly bell-shaped with a mean of $43 and standard deviation of $15. Construct a confidence interval at a 95% confidence level.
Give your answers to one decimal place.

Answer by Theo(13342)

sample size is 11.
sample mean is 43 dollars.
standard deviation is 15 dollars.
because you are looking for the mean of a sample of specified size, use the standard error rather than the standard deviation.
standard error = standard deviation divided by square root of sample size = 15 / sqrt(11) = 4.522670169.
at 95% two tailed confidence level, the critical t-score with 10 degrees of freedom is equal to plus or minus 2.228138842.
use the t-score formula to find the equivalent raw score.
on the high side, t = (x - m) / s becomes 2.228138842 = (x - 43) / 4.522670169.
solve for x to get x = 2.228138842 * 4.522670169 + 43 = 53.07713707.
on the low side, t = (x - m) / s becomes -2.228138842 = (x - 43) / 4.522670169.
solve for x to get x = -2.228138842 * 4.522670169 + 43 = 32.92286293.
your 95% confidence interval is from 32.92286293 to 53.07713707.
in the t-score formula for this problem, .....
t is the critical t-score
x is the critical raw score
m is the mean
s is the standrd error.